### v1

# Language Model Council: Benchmarking Foundation Models on Highly Subjective Tasks by Consensus

Abstract
The rapid advancement of Large Language Models (LLMs) necessitates robust and challenging benchmarks. Leaderboards like Chatbot Arena rank LLMs based on how well their responses align with human preferences. However, many tasks such as those related to emotional intelligence, creative writing, or persuasiveness, are highly subjective and often lack majoritarian human agreement. Judges may have irreconcilable disagreements about what constitutes a better response. To address the challenge of ranking LLMs on highly subjective tasks, we propose a novel benchmarking framework, the Language Model Council (LMC). The LMC operates through a democratic process to: 1) formulate a test set through equal participation, 2) administer the test among council members, and 3) evaluate responses as a collective jury. We deploy a council of 20 newest LLMs on an open-ended emotional intelligence task: responding to interpersonal dilemmas. Our results show that the LMC produces rankings that are more separable, robust, and less biased than those from any individual LLM judge, and is more consistent with a human-established leaderboard compared to other benchmarks.
1 Introduction
We are experiencing a Cambrian Explosion of Large Language Models (LLMs) that exhibit remarkable and wide-ranging abilities. As LLMs have advanced, evaluating their quality has become increasingly challenging. Publishers of new models often claim superiority based on various benchmarks, citing their rankings on leaderboards like [6, 10, 15, 25, 28].
Conventional LLM evaluations use closed-ended questions (e.g., multi-choice questions like MMLU [20] or TruthfulQA [26]) that can be automatically verified. However, the static nature of benchmarks risks contamination, where models may have been inadvertently exposed to elements of the test datasets during training, thereby skewing evaluation results [31].
Figure 1: Overview of benchmarking by Language Model Council. Using the same LLMs for test set formulation, task completion, and judging offers an equitable way to achieve a ranking by consensus on highly subjective tasks.
Arena-based approaches such as [44, 7, 10, 43, 15, 24] aim to address this limitation by evaluating LLMs on open-ended questions. These methods typically involve two models competing in various tasks to evaluate their abilities. Examples of these tasks include responding to open-ended questions, completing specific tasks, playing skill-based games, or having multi-turn conversations. The outcomes of these battles serve as indicators of each model’s competency and can be determined objectively (NegotiationArena [7]), by a strong judge model like GPT-4 (AlpacaEval [15], Chatbot Arena Hard [24]), by a committee of LLM judges (PoLL [36], Auto Arena [43]), or by real humans (MT-Bench and Chatbot Arena [24]).
But how consistent are human ratings to begin with? Substantial disagreements are common in labeling online comment toxicity, news misinformation, and medical diagnosis, with up to one-third of expert annotators disagreeing on an average example, even after accounting for label noise [18]. In a recent study on persuasiveness, humans expressed different experiences after reading the same argument 89% of the time [16]. On MT-Bench, human ratings agree only 38% of the time on a per-item basis, with 0% unanimity among expert raters [44].
There is growing interest in establishing benchmarks for emotional intelligence (EI) in Large Language Models (LLMs) [28, 37, 32]. However, assessing EI is highly subjective, and the relationship between EI and IQ remains contentious, with studies showing positive, negative, or no correlation [27, 17]. On EmoBench, a handcrafted test for emotional intelligence, only 32.5% of questions receive unanimous agreement, despite its simple, multiple-choice format [32].
To tackle the challenge of evaluating LLMs on highly subjective tasks, we introduce the Language Model Council (LMC). The LMC consists of three stages. First, a test set is formulated with equal participation from all council members. Second, the test set is administered to all council members to complete. Lastly, the responses are evaluated by the council as a collective jury.
Our contributions are as follows:
  • 1.
    We propose the LMC as a flexible evaluation framework for ranking LLM agents on subjective tasks in a democratic fashion. We rank 20 state-of-the-art LLMs based on an open-ended emotional intelligence task that is more separable, robust, and less biased than using any individual LLM judge. Qwen-1.5-110B emerges as the leader, surpassing GPT-4o in second place.
  • 2.
    We define and quantify key measures of LLM judging quality in an ensemble setting such as consistency, agreement, separability, affinity, and bias. We identify claude-3-opus and mistral-large as the most effective judges.
  • 3.
    We run a human study to determine a parallel leaderboard on this EI task for a subset of 9 LLMs. We find that the LMC’s ranking is significantly more correlated with the human-established rankings compared to other benchmarks.
  • 4.
    We use Monte-Carlo simulations to measure ranking stability and robustness to adversarial judges. We discuss the trade-offs between efficiency and inclusivity in larger LLM evaluation ensembles and the value of the incremental LLM judge.
We release all data, code, and leaderboard at https://llm-council.com.
2 Related Work
Disagreement and dissenting voices amongst humans and LLMs. Data perspectivism [5, 30, 33] posits that differing opinions from human annotators on subjective phenomena can all be correct, reflecting their diverse cultural and demographic backgrounds, and should thus be considered true in the gold standard. In earlier language models, disagreements with human annotations were often seen as errors, hallucinations, or incompetencies. However, as LLMs exhibit more sophisticated abilities, these differences in judgment may be increasingly interpreted as valid dissenting perspectives. Notably, when humans disagreed with GPT-4, they deemed GPT-4’s judgments reasonable in 75% of cases and were willing to change their choices in 34% of cases [44]. LLMs, influenced by human data selection and algorithm design, exhibit distinct personalities [35] and biases [2, 13, 22, 29].
Using LLMs for synthetic data. LLMs can be used as a source of synthetic data for training or evaluation [38, 19, 41, 4]. By utilizing LLM(s) to generate questions, these methods can effectively mitigate the contamination concerns on static datasets and sometimes create boosts in test set variety.
LLM evaluation ensembles. Recent works have proposed using multiple agents for collaborative evaluation, mimicking a peer review process. Language-Model-as-an-Examiner [4] employs LLMs to interact with candidates through follow-up fact-oriented queries in the knowledge domain. Auto Arena [43] utilizes an LLM committee to judge competitive multi-turn interactions between LLMs. PRD [23] allows LLMs to discuss evaluations, assigning higher voting weights based on ability. PRE [11] selects a small group of reviewers to produce individual evaluations, then aggregates these evaluations through a chair. DRPE [42] uses multi-roleplayer prompting to simulate different roles with the same base model and integrates multiple outputs as votes for the final results. PoLL [36] enhances cost-effectiveness by replacing a single large judge with multiple smaller judges.
We build upon existing research by: 1) focusing on a highly subjective task where human agreement is inherently low, such as emotional intelligence, 2) using the same set of LLMs end-to-end for test set formulation, task completion, and response judging to achieve a more equitable consensus, akin to democratic social structures, and 3) engaging a large ensemble of diverse LLMs to study judging dynamics more richly. To our knowledge, this is the largest panel of LLM judges studied to date.
3 Case Study: Using the LMC to Rank LLMs on Emotional Intelligence
The Language Model Council framework consists of three stages: test set formulation, response gathering, and collective judging (Figure 1). These three stages are run sequentially and fully automated with LLM-based agents. All prompts are included in Appendix F.
3.1 Council member selection
Our selection of council member models was guided by several key considerations, including their widespread adoption within the AI community, availability of technical reports, well-supported API access, and performance on benchmarks like MMLU [20] and LMSYS [10]. We ensure a broad variety of LLMs by including models from eight different organizations across four countries, with a mix of open and closed-source models, small and large models (Figure 8).
3.2 Test set formulation and response gathering
To create a compelling open-ended test set for Emotional Application (EA), we build upon the EmoBench dataset, a hand-crafted, theory-based dataset designed for this purpose [32]. EmoBench consists of 200 emotionally balanced, handcrafted questions, e.g., “Sarah found out that her younger brother is being bullied at school, but he begged her not to tell their parents.” We use the Council to transform the concise, closed-ended scenarios in EmoBench into richly described, open-ended dilemmas in the first person. See Figure 21 for the exact prompt used. Due to budget constraints, each of the 20 council members expands 5 scenarios, resulting in a dataset of 100 dilemmas, similar in scale to MT-Bench (80 questions). We manually review all expansions for EA suitability.
Relying on a single LLM to generate the entire test set, even a top performer like GPT-4o (#1 on Chatbot Arena [^1] https://chat.lmsys.org/), may introduce bias and limit perspective diversity. In a survey of 10 human respondents, 51% preferred expansions not written by GPT-4o to be included in an EA test, highlighting the subjectivity of test set construction and the need for diverse input. Inclusively constructing test sets mitigates the risk of any single LLM’s generative idiosyncrasies [1] from dominating.
Another option for inclusive test set formation is having multiple council members propose expansions for each scenario and then selecting the best through voting. This is viable but resource-intensive. Given the high quality of expansions, the incremental value of this approach is small compared to simply expanding the test set with new scenarios.
Once all 100 scenarios have been generated, each council member responds to each, resulting in 2,000 responses. In the prompt (Figure 22), we request adherence to a 250-word limit to mitigate length bias [14]. Responses exceeding this limit are truncated to the last sentence within the limit.
3.3 LLMs-as-a-jury evaluation settings
LLM rankings are based on expected win rates derived from an ELO scoring system [3, 8]. Similar to the Chatbot Arena score calculation procedure [10], we compute the Bradley-Terry (BT) coefficients [9] for better statistical estimation.
4-point preference scale. Based on the results of a calibration exercise (Appendix B), we query all LLMs with a temperature of 0 and use granular comparison options without ties (A>>B, A>B, B>A, B>>A) in the prompt. We use Chain-of-Thought (CoT) prompting [40] to generate answers before giving judgments.
Choice of reference respondent. Following Arena Hard [24], we use a common reference model for all pairwise battles. A dry run with 5% of the data compared GPT-4-0613 and Qwen-1.5-32B as reference models. For the main experiment, we proceed with Qwen-1.5-32B as the reference model since the battle results with Qwen-1.5-32B were more varied and resulted in significantly more separable ELO scores.
Dual-sided battles. To mitigate position bias, we adopt a two-game setup, swapping model positions per query, resulting in $100*2=200$ judgments per model per judge. Following the implementation of Bradley-Terry coefficient calculations in the original codebase [^2] https://github.com/lm-sys/arena-hard-auto, inconsistent results after swapping are treated as ties and strong votes are counted as 3 separate wins.
Voting aggregation functions. We consider 3 different voting functions for aggregating scores across multiple LLM judges: no aggregation (each battle judgment is equally considered); majority vote (for a given battle, we use the mode of the votes from all council members), and mean pooling (ratings are mapped to a 4-point scale (A>>B: 2, A>B: 1, B>>A: -1, B>>A: -2) and the mean of which is rounded to the nearest whole value).
3.4 Characterizing LLM judges
We leverage the many-to-many interactions between LLMs to quantify key measures of LLM judging quality in an ensemble setting.
Separability quantifies how well models can be distinguished using confidence intervals (via bootstrapping) [24]. It is measured by the percentage of model pairs with non-overlapping score confidence intervals. Higher separability indicates better model differentiation by the LLM judge. Conviction is the percentage of strong votes (e.g., A>>B or B>>A).
Consistency measures how often a judge gives consistent results when the order of two assistants is swapped. Position bias refers to an LLM favoring certain positions in pairwise comparisons [44]. Inconsistent votes are mapped to favoring either the first or second position, described in Table 10.
Agreement is measured using Cohen’s Kappa [12] between two judges’ ratings. To avoid penalizing minor order-consistent differences, granular ratings are first mapped to coarse sidewise buckets. Contrarianism is defined as the percentage of disagreement between an LLM and the Council’s majority decision.
Affinity between a judge and respondent equals the win rate the respondent model receives under the judge’s jurisdiction. Self-enhancement bias is the difference between a model’s affinity to itself and the council’s overall affinity to it. Polarization is the range of the highest and lowest assigned scores. Length bias is the $R^{2}$ of a linear regression model predicting score from average response length.
4 Results and Findings
As a Respondent As a Judge
LLM Rank Council EI Score Avg. response length Separability Consistency
qwen1.5-110B-Chat 1 65.6 (-1.2, 1.8) 233 62.1% 67.6%
gpt-4o-2024-05-13 2 59.2 (-1.2, 1.7) 224 60.5% 50.8%
gpt-4-turbo-2024-04-09 3 57.5 (-1.2, 1.7) 221 57.9% 38.5%
gemini-1.0-pro 4 50.6 (-1.2, 1.5) 228 30.5% 34.8%
claude-3-opus 5 50.1 (-1.5, 1.4) 228 72.6% 74.6%
qwen1.5-32B-Chat 6 50.0 (0.0, 0.0) 236 25.3% 23.5%
qwen1.5-72B-Chat 7 48.7 (-1.4, 1.6) 236 37.9% 26.9%
llama-3-70b-chat 8 45.1 (-1.5, 1.4) 224 64.2% 51.1%
claude-3-sonnet 9 42.5 (-1.5, 1.6) 226 52.1% 39.7%
dbrx-instruct 10 38.8 (-1.5, 1.9) 233 50.5% 44.2%
claude-3-haiku 11 38.6 (-1.7, 2.2) 234 45.3% 44.2%
command-r-plus 12 35.6 (-1.7, 1.7) 222 61.1% 52.9%
command-r 13 34.7 (-1.7, 1.5) 227 45.8% 54.5%
mixtral-8x7b 14 34.4 (-1.4, 1.5) 233 56.8% 58.6%
mistral-large 15 33.9 (-1.5, 1.3) 208 73.7% 72.5%
llama-3-8b-chat 16 30.0 (-1.4, 1.4) 207 31.1% 26.1%
mistral-medium 17 29.3 (-1.6, 1.5) 185 57.9% 59.0%
gpt-4-0613 18 26.9 (-1.4, 1.4) 173 64.7% 53.6%
gpt-3.5-turbo-0125 19 18.2 (-1.1, 1.1) 187 55.8% 57.7%
gemini-1.5-pro 20 11.6 (-0.9, 0.8) 115 60.0% 52.3%
Average Judge 53.3% 49.2%
council (by majority vote) 73.7% 75.3%
council (by mean pooling) 74.7% 68.5%
council (no aggregation) 90.5% 52.3%
Table 1: The LMC promotes equal participation as respondents and judges. The Council EI rank and scores are derived from the “council (no aggregation) setting,” where ratings from all LLMs are tallied verbatim, without aggregation or modification. Under various aggregation algorithms, the council is more separable and more consistent than individual LLM judges.
4.1 On Respondents
Table 1 shows the main results of our Council-based EI study. In stark contrast to other leaderboards, Qwen-1.5-110B (#20 on Chatbot Arena) attains the highest ELO score on our EI task, followed by GPT-4o (#1 on Chatbot Arena) in second place. This may be due to the use of its predecessor, Qwen-1.5-32B, as the only common reference model, giving Qwen-1.5-110B a family-specific advantage. A qualitative analysis in Section 5.2 investigates this further. Within other model families, larger models generally rank higher than their smaller, older counterparts, with Gemini-1.0-pro (ranked 4th) versus Gemini-1.5-pro (ranked last) being a notable exception.
Losers ignore word limits. Despite a suggested 250-word limit, a handful of models consistently generate much shorter responses even with ample max_new_tokens. The top 14 models used most of the allowed word length (220+ words), while all models in the bottom 4 averaged less than 200. Notably, Gemini-1.5-pro, despite being the successor to Gemini-1.0-pro, places last with responses averaging 115 words, less than half of what was allowed. Adherence to suggested word limits may be an area of improvement for LLMs and something to consider for future length-conscious benchmarks.
4.2 On Judges
Models with lower invariability during calibration (Appendix B) were also less consistent in the main experiment, indicating that calibration can be used to anticipate a judge’s performance.
Variable Correlation
Council EI Score -0.04
Consistency 0.74
Conviction -0.44
Contrarianism -0.78
Polarization 0.86
Table 2: Spearman correlation between EI score and key judging qualities to separability.
Consistency strongly correlates with separability (Table 2), likely due to reduced noise in ratings resulting in tighter confidence intervals. Polarization also shows strong correlation, indicating respondents are evenly clustered with moderate outliers. qwen1.5-72B-Chat has the highest contrarianism score but only 38.9% separability. Notably, claude-3-opus achieves the 2nd highest separability despite expressing strong preferences only 0.1% of the time. Conversely, llama-3-8b-chat has the highest conviction, but only 31.1% separability. These findings suggest that the quality of votes is more critical for separability than the strength or incongruity of opinion. Finally, the Council’s EI score is weakly correlated with key judging qualities, indicating that effective judging and task completion are distinct skills.
Table 12 shows that 12 out of 20 LLMs show positive self-enhancement bias, with llama-3-* models exhibiting the highest levels (+0.11 for 70b, +0.21 for 8b). Interestingly, six out of 20 LLMs exhibit negative self-bias. The overall ranking remains similar when self-graded battles are excluded, confirming the effectiveness of using an ensemble of LLMs to mitigate self-bias. While length bias is high, excluding the bottom four models, which were significantly under-length, makes length bias insignificant (Table 3), suggesting that much shorter responses are systematically not preferred.
Models with $>$200 words All models
All votes Consistent votes All votes Consistent votes
Average judge 0.158 0.143 0.502 0.319
Council (majority vote) 0.116 0.129 0.365 0.354
Council (mean pooling) 0.112 0.139 0.592 0.389
Council (no aggregation) 0.125 0.106 0.545 0.347
Table 3: Length bias with and without models with responses $<$200 words.
4.3 Comparison with Human Performance
Human GPT-4o C-A C-M)
Human 51.9% 51.4% 52.2% 54.2%
GPT-4o 51.4% 60.2% 78.6%
C-A 52.3% 60.2% 56.4% 67.4%
C-M 54.2% 56.4% 67.4%
Table 4: Agreement between humans and the LMC on the LMC’s EI task. "C-A" denotes a body of 20 individual LLMs while "C-M" is the Council with majority aggregation.
Figure 2: LLM rankings from different systems.
To validate the Council’s evaluations, we conduct a human study mirroring that with the Council: human raters are asked to choose the best of two responses to a given dilemma. We also invite participants to justify their choice based on perceived EI and empathy (adapted from the PETS questionnaire [34]), actionableness, clarity and conciseness. In addition, they have the option to provide a free-text explanation for their choice.
We select nine LLM council members from our pool of 20 for this study, ensuring variety in terms of model size, open- versus closed-source, and company origin. They are claude-3-opus, Qwen1.5-110B, gpt-4o, Llama-3-70b, claude-3-haiku, Llama-3-8b, Qwen1.5-32B-Chat, and gpt-4-0613.
We recruit participants via crowdsourcing on Prolific [^3] https://www.prolific.com/. We randomly sample 120 dilemmas-response triples from the LLM generation. We recruited a total of 102 participants. Each dilemma pair and response were rated by 11 participants on average, with a total of 1343 ratings. Detailed information on our recruitment process, quality control, and participant demographics is provided in Appendix D.
We follow the same methodology described in Section 3.3. As shown in Figure 2, the relative ranking of the systems is consistent between the humans and the LLM council, with small changes. The three top-performing models, as well as the bottom ones, emphasizing a consensus on the ranking of the best and worse models. Compared to the human agreement of 65% on MT-Bench [^4] Agreement between two judges is defined as the probability of randomly selected individuals (but not identical) of each type agreeing on a randomly selected battle., Figure 4 confirms that our EI task is more subjective, with a lower human agreement of 51.9%. Still, the level of agreement between humans and the Council with and without aggregation is roughly the same as the agreement between humans.
4.4 Comparison to Other Leaderboards
Figure 3: Spearman correlation analysis on evaluation benchmarks for 9 LLMs.
Since our case study aims to rank models on a narrow but subjective EI task, high correlation with Chatbot Arena, which measures general capability, is neither intended nor a desired outcome of our project. Instead, Figure 3 shows that the Council successfully exhibits high correlation with the human study, despite the highly subjective nature of our EI task.
5 Discussion
5.1 Can Judging Consistency be Improved with Oligarchical Councils?
Table 1 shows a wide variety of judging qualities within the council, especially in consistency and separability. In the two-game setup, we collect ratings for models in both positions of pairwise comparisons, which allows us to identify and potentially remove inconsistent ratings before ELO scoring. To explore the dynamics of voting aggregation and the exclusion of inconsistent votes, we formed three oligarchical councils with subsets of LLMs: Flagships, the largest LLM from each organization; Smalls: the smallest LLMs from each organization; and Top-4, the top 4 LLMs according to Chatbot Arena. See Table 9 for a detailed list. The ratings from the main experiment are reused to assess what “would have been” if the sub-council was used instead of the full council.
Figure 4: Separability scores achieved by different council compositions and aggregation methods.
Across all council compositions, the separability of majority aggregation is always higher than the average judge and often higher than the best judge. The council without aggregation achieves the highest separability, probably due to more ratings, which generally means tighter confidence intervals.
However, the smallest council, even with more votes, cannot reach the separability level of the best council with just four judges. Moreover, using only consistent votes means that sometimes  50% of ratings are filtered out, yet separability scores increase.
The council of the smallest models has the highest rate of inconsistent votes and the largest boost in separability when these votes are removed. In all configurations, the top 2 ranked models are always Qwen1.5-110B and GPT-4o, respectively.
To achieve the highest separability and sharpest confidence intervals, especially in low data scenarios, we recommend using no aggregation in LLM evaluation ensembles.
Council Composition Separability Conviction Consistency Polarization Length bias
all 0.92 (+0.01) 0.05 (+0.04) 1.0 (+0.48) 0.81 (+0.27) 0.35 (-0.19)
flagships 0.90 (+0.03) 0.04 (+0.03) 1.0 (+0.48) 0.85 (+0.22) 0.29 (-0.17)
smalls 0.81 (+0.10) 0.11 (-0.21) 1.0 (+0.74) 0.73 (+0.26) 0.44 (-0.25)
top-4 0.86 (+0.07) 0.04 (+0.03) 1.0 (+0.48) 0.84 (+0.24) 0.25 (-0.13)
Table 5: Changes to key judging qualities when only using consistent votes in the (no aggregation) council.
5.2 Qualitative Analysis: What Makes a Response Preferred Over Another?
Several arena-based benchmarks (ours included) have demonstrated that a clear ranking among LLMs can be established, but there is a dearth of analysis and understanding as to why the rankings are the way they are. For example, platforms like Chatbot Arena do not clarify how factors like feel and style are weighed against correctness [39], and CoT explanations from MT-Bench remain unanalyzed.
We aim to better understand the qualitative aspects of what makes a response to an emotional interpersonal conflict more desirable to better inform how to improve future models. First, we manually examine a random sample of 50 explanations, identifying 38 different reasons for preferences (e.g., “more practical”). The full list is in Appendix E. Next, we used a strong LLM (GPT-4o) to map a larger sample of 1K explanations to these predefined reasons (prompt in Figure 26). The 1K sample includes ratings from all 20 LLM judges. Detailed reason citation frequencies are listed in Figure 19.
We find that LLM judge-provided ratings are almost always based on multiple indicators (4.5 ± 2.4 on average). "More actionable" is the most cited reason, which aligns with the action-oriented framing of our emotional intelligence test. "Structure," "clarity," and "specificity" dominate the top 10 reasons. "More gentle" and "more soft" are cited least, contrasting with "more practical" (#11) and "more authentic" (#12). Longer responses ("more comprehensive” #2, "more detailed" #3) are preferred over brevity ("less verbose" #9).
We also examine feedback from the human study: we find that users generally find that the best responses display emotional intelligence (60.9%), are actionable (55.1%) and clear (52.9%). In contrast, participants reported the best response is concise only 15.9% of the time, suggesting length is less of a determining factor for humans. Moreover, we find little support for empathy: the participants did not find any of the statements in the PETS questionnaire to ring any truer for the winning response. Participants who provided verbal feedback emphasized specificity to the situation, clear examples of how to proceed, and a tone that was not too formal. Full details in Appendix D.
Overall, feedback from human participants and automatically extracted reasons from LLM judge explanations share consistent themes: longer responses that are clear, detailed, and actionable are better when responding to emotional interpersonal conflicts.
Reason 1 Reason 2 Correlation
less verbose more succinct 0.650
better structured more structured 0.584
easier to follow better structured 0.520
easier to follow more structured 0.468
less verbose more direct 0.450
more understanding more empathetic 0.418
more clear better structured 0.387
more direct more succinct 0.349
easier to follow more clear 0.348
more gentle more soft 0.337
Table 6: Top 10 positive correlations.
Reason 1 Reason 2 Correlation
more comprehensive less verbose -0.276
less verbose more detailed -0.227
more comprehensive more direct -0.202
more comprehensive more succinct -0.197
more detailed more succinct -0.196
more comprehensive more focused -0.161
more detailed more direct -0.148
more suggestions … less verbose -0.144
more understanding more actionable -0.139
less verbose more nuanced -0.135
Table 7: Top 10 negative correlations.
5.3 Jury Ablation: What is the Value of the Incremental Judge?
If there are $n$ test prompts and $m$ council members, the cost of evaluating a single LLM is $O(nm)$, or $O(nm^{2})$ for $m$ LLM judges. Unsurprisingly, full participation is costly – as $m$ increases, evaluation costs rise significantly compared to using a single LLM judge.
Councils with full participation achieve the highest separability scores. However, strong oligarchical councils also show competitive separability, suggesting that full participation may not be necessary for sufficiently good rankings. Conversely, councils of smaller models exhibit worse judging quality, emphasizing the need for thoughtful council composition.
To explore the trade-off between efficiency and inclusivity in determining council size and composition, we use Monte Carlo methods to simulate councils of different sizes and compositions, assessing what the rankings would be with different subsets of LLM juries.
We quantify stability by measuring the Mean Variance of the Rank (MVR), defined as the variance of a respondent’s rank averaged over all respondents and random trials. An MVR of 1 means a respondent’s rank is expected to change by 1 position in a new trial. Lower MVR indicates more stable rankings, with MVR of 0 signifying perfect stability. To assess robustness, we observe the impact of adding adversarial judges who vote randomly. Results for 100 random samples for increasing council sizes are shown in Figure 5.
Figure 5: Monte Carlo jury ablation measuring mean variance of the rank (MVR) (left) and separability (right) using majority-aggregated juries. The x-axis represents the number of non-adversarial judges in each jury. Each data point represents the score over 100 randomly sampled jury compositions. The 95% confidence intervals are shown for MVR. For separability, random trials bootstrap confidence intervals directly.
Larger juries have lower MVR, both with and without adversarial judges. Adversarial judges negatively impact both MVR and separability. Larger councils demonstrate consistent robustness to adversarial judges in both MVR and separability, though with diminishing marginal returns, flattening around councils of size 12.
6 Limitations
We only studied single-turn arguments. Our study evaluates EI based on exposure to single, self-contained interactions. In many contexts, EI is effectively applied through iterative discussions. A more interactive setup with dynamic exchanges could create stronger tests for EI.
Cultural and linguistic context. Our study focuses on English articles and English speakers. Further research would be needed to determine the broader applicability of our results to other languages.
Resource allocation can be significant for large councils. While the assumption of equal weight for all council opinions guides this work, effective subsets and variations in judging proficiency highlight complexities. Balancing example quantity and council size requires careful consideration.
Reproducibility. Closed-weight models like GPT-4 pose a reproducibility challenge due to undisclosed updates, causing responses to vary despite fixed settings. To ensure reproducibility, we provide details like the source and hyperparameters used.
Emotional responses are personal, influenced by experiences and social factors. The effectiveness of LLM’s response is best judged by the person experiencing the emotion. Our quantitative evaluation may not fully capture the essence of a first-person judgment of conversational responses.
7 Conclusion
In this paper, we propose the LMC as a flexible evaluation framework for ranking LLM agents in a fully democratic fashion. The framework is flexible, easily extendable, and its decentralized design is inherently robust to adversarial participants. An in-depth case study applying the Council to a highly subjective emotional intelligence task with 20 state-of-the-art LLMs shows that Council, in a fully decentralized manner, produces highly separable rankings that correlate significantly more with human-established rankings than other benchmarks. Altogether, the LMC establishes a foundation for a consensus-based LLM evaluation framework for highly subjective tasks.
Acknowledgements
Justin led the research. Amanda and Flor led the human study. We thank Sahand Sabour for creating EmoBench and insightful discussions about emotionally rich synthetic data. We thank Alex Tamkin for proposing the idea to measure the value of the incremental judge. We thank Mitchell Gordon for his suggestion to qualitatively analyze why certain responses are preferred. We thank David So for his ideas on calibration, repeatability, and model affinity. We thank Federico Bianchi and Sam Paech for their idea of measuring oligarchical councils and for their feedback on prompt design. Sam Paech provided insightful discussions on separability, voting aggregation, length bias, and the relationship to other leaderboards, and he advised us throughout the project. Finally, we thank Predibase for their support and funding this research.
Appendix A Language Model Council Configurations
Country Organization LLM Release Date Chat Arena Elo MMLU (5-shot) Size License
United States Open AI gpt-4o-2024-05-13 05/24 1287 88.7 [Uncaptioned image] Proprietary
United States Open AI gpt-4-turbo-04-09 04/24 1256 [Uncaptioned image] [Uncaptioned image] Proprietary
United States Open AI gpt-4-0613 06/23 1246 86.4 [Uncaptioned image] Proprietary
United States Open AI gpt-3.5-turbo-0125 01/24 1102 70.0 [Uncaptioned image] Proprietary
France Mistral mistral-large-latest 02/24 1156 81.2 [Uncaptioned image] Proprietary
France Mistral open-mixtral-8x22b 04/24 1146 77.8 176 B Apache 2.0
France Mistral open-mixtral-8x7b 12/23 1114 70.6 56 B Apache 2.0
United States Meta llama-3-70b-chat-hf 04/24 1208 82.0 70 B Llama 3 Community
United States Meta llama-3-8b-chat-hf 04/24 1153 68.4 8 B Llama 3 Community
United States Google gemini-1.5-pro-preview-0409 05/24 1268 81.9 [Uncaptioned image] Proprietary
United States Google gemini-1.0-pro 04/24 1208 71.8 [Uncaptioned image] Proprietary
United States Databricks dbrx 03/24 1103 73.7 132 B DBRX LICENSE
Canada Cohere command-r-plus 04/24 1189 75.7 104 B CC-BY-NC-4.0
Canada Cohere command-r 04/24 1147 68.2 35 B CC-BY-NC-4.0
United States Anthropic claude-3-opus-20240229 03/24 1248 86.8 [Uncaptioned image] Proprietary
United States Anthropic claude-3-sonnet-20240229 03/24 1201 79.0 [Uncaptioned image] Proprietary
United States Anthropic claude-3-haiku-20240307 03/24 1178 75.2 [Uncaptioned image] Proprietary
China Alibaba qwen1.5-110B-chat 02/24 1164 80.2 100 B Qianwen LICENSE
China Alibaba qwen1.5-72B-chat 02/24 1152 77.4 72 B Qianwen LICENSE
China Alibaba qwen1.5-32B-chat 02/24 1126 74.3 32 B Qianwen LICENSE
Table 8: 20 council members used for experiments in this work. We include models from eight different organizations across four countries, with a mix of open and closed-source models, small and large models. To our knowledge, this is the largest panel of LLM judges studied to date.
LLM All Flagships Smalls Top-4
gpt-4o-2024-05-13 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
gpt-4-turbo-04-09 [Uncaptioned image] [Uncaptioned image]
gpt-4-0613 [Uncaptioned image]
gpt-3.5-turbo-0125 [Uncaptioned image]
mistral-large-latest [Uncaptioned image] [Uncaptioned image]
open-mixtral-8x22b [Uncaptioned image]
open-mixtral-8x7b [Uncaptioned image] [Uncaptioned image]
llama-3-70b-chat-hf [Uncaptioned image] [Uncaptioned image]
llama-3-8b-chat-hf [Uncaptioned image] [Uncaptioned image]
gemini-1.5-pro-preview-0409 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
gemini-1.0-pro [Uncaptioned image] [Uncaptioned image]
dbrx [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
command-r-plus [Uncaptioned image] [Uncaptioned image]
command-r [Uncaptioned image] [Uncaptioned image]
claude-3-opus-20240229 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
claude-3-sonnet-20240229 [Uncaptioned image]
claude-3-haiku-20240307 [Uncaptioned image] [Uncaptioned image]
qwen1.5-110B-chat [Uncaptioned image] [Uncaptioned image]
qwen1.5-72B-chat [Uncaptioned image]
qwen1.5-32B-chat [Uncaptioned image] [Uncaptioned image]
Table 9: Additional council variations consisting of a hand-picked subset of LLMs.
Appendix B LLM Judge Calibration
Figure 6: The scenario, synthetic expansion, and responses used for pairwise comparison calibration. Three possible responses are evaluated: one from Claude Opus, one from GPT-4o, and a generic response: “I’m sorry it sounds like you are going through a rough time. I wish you the best.”
To understand the consistency and variability of LLM model judges under different temperatures and pairwise comparison options, we collect pairwise preference ratings on three responses to the same interpersonal conflict. Two responses are competitive, and one is intentionally generic to serve as a ranking baseline (Figure 6). We measure:
Invariability:
How reliably does the model give the same answer every time with the same pair in the same order?
Let:
  • $P$ be the set of all pairs of responses.
  • $R_{i,j}$ be the result of the $j$-th repetition of the pairwise comparison of the $i$-th pair $(x_{i},y_{i})$ in the same order.
  • $n$ be the number of repetitions.
For each pair $(x_{i},y_{i})$, we perform $n$ comparisons, resulting in a set of results $\{R_{i,1},R_{i,2},\ldots,R_{i,n}\}$.
Define the mode of the set $\{R_{i,1},R_{i,2},\ldots,R_{i,n}\}$ as $\text{mode}(R_{i})$.
The frequency of the mode for the $i$-th pair is given by:
$f_{i}=\frac{\sum_{j=1}^{n}\mathbb{I}(R_{i,j}=\text{mode}(R_{i}))}{n}$
where $\mathbb{I}$ is the indicator function, which is 1 if the condition inside is true, and 0 otherwise.
The invariability is then defined as the average of $f_{i}$ over all pairs in $P$:
$invariability=\frac{1}{|P|}\sum_{i\in P}f_{i}$
Consistency:
Does the model give a consistent answer when the order of the same pair of respondents is flipped?
To measure consistency over several repetitions of the same items in both orders, we take the percentage of consistent couplets over all possible rating couplets.
Let:
  • $P$ be the set of all pairs of responses.
  • $R_{i,j}$ be the result of the $j$-th repetition of the pairwise comparison of the $i$-th pair $(x_{i},y_{i})$ in the same order.
  • $R_{i^{\prime},j}$ be the result of the $j$-th repetition of the pairwise comparison of the $i$-th pair $(y_{i},x_{i})$ in swapped order.
  • $n$ be the number of repetitions.
For each pair $(x_{i},y_{i})$, we perform $n$ comparisons in both the original and swapped orders, resulting in two sets of results: $\{R_{i,1},R_{i,2},\ldots,R_{i,n}\}$ and $\{R_{i^{\prime},1},R_{i^{\prime},2},\ldots,R_{i^{\prime},n}\}$.
[…]
Consistency is then defined as the average consistency over all pairs $(i,j)\in P$ and repetitions:
$consistency=\frac{1}{|P|\cdot n^{2}}\sum_{i\in P}\sum_{j=1}^{n}\sum_{k=1}^{n}\text{are\_consistent}(R_{i,j},R_{i^{\prime},k})$
This is equivalent to the percentage of consistent couplets over all possible rating couplets.
To assess invariability, each LLM judge is prompted 5 times with the original pairwise comparison prompt (Figure 24 and 5 times with a trivially reworded version of the prompt  [^5] Trivial rewording involves changing the first sentence of the judging prompt (Figure 24) to: ”This person is experiencing an emotional dilemma and is seeking guidance and help.”. To assess consistency, each pair of responses is evaluated in both orders, also for 10 repetitions for each.
In our calibration, there are 10 reps of each pair of responses in both positions, so there are $10*10=100$ instances of swapped-position rating couplets. The are_consistent function for consistency metrics is based on the mapping defined in Table 10.
Rating Order-swapped rating Consistent Inconsistent Biased towards first Biased towards second
A>>B or A>B A>>B or A>B FALSE TRUE TRUE FALSE
A>>B or A>B B>>A or B>A TRUE FALSE FALSE FALSE
A>>B or A>B Asimilar-$\sim$=B FALSE TRUE TRUE FALSE
B>>A or B>A A>>B or A>B TRUE FALSE FALSE FALSE
B>>A or B>A B>>A or B>A FALSE TRUE FALSE TRUE
B>>A or B>A Asimilar-$\sim$=B FALSE TRUE FALSE TRUE
Asimilar-$\sim$=B A>>B or A>B FALSE TRUE TRUE FALSE
Asimilar-$\sim$=B B>>A or B>A FALSE TRUE FALSE TRUE
Asimilar-$\sim$=B Asimilar-$\sim$=B TRUE FALSE FALSE FALSE
Table 10: Reference table for categorizing a couplet of order-swapped ratings of the same set of items, (A, B) vs. (B, A). We do not penalize consistency as long as the overall side of the preference is consistent. All inconsistent votes are either biased towards the first or second position.
We test 3 different temperatures ($0.0$, $0.5$, $1.0$) and $4$ different sets of pairwise comparison options:
  • Coarse preferences with tie option (A>B, B>A, Asimilar-$\sim$=B)
  • Coarse preferences without tie option (A>B, B>A)
  • Granular preferences with tie option (A>>B, A>B, B>A, B>> A, Asimilar-$\sim$=B)
  • Granular preferences without tie option (A>>B, A>B, B>A, B>>A)
Figure 7: Calibration scores for invariability (left) and consistency (right), averaged over 20 LLMs and 10 repetitions for each under different pairwise comparison options.
Somewhat surprisingly, several models do not produce fully invariant ratings, even with $temperature=0$. However, using a lower temperature increases consistency. Excluding the tie option slightly improves invariability and consistency at some temperatures, with no negative impact at $temperature=0$. Table/Figure [table ref] lists detailed calibration results for individual LLMs with granular pairwise comparison options without ties and with $temperature=0$.
Our calibration study concludes with the decision to use granular comparison options without a tie to "force" judges to choose a side, thereby better distinguishing models, and with $temperature=0$.
LLM Invariability Conviction Consistency Position bias Position bias
(strong votes) (first) (second)
claude-3-haiku 100.0% 50.0% 50.0% 50.0% 50.0%
claude-3-opus 100.0% 50.0% 100.0% 0.0% 0.0%
claude-3-sonnet 100.0% 50.0% 100.0% 0.0% 0.0%
command-r 100.0% 50.0% 50.0% 50.0% 50.0%
command-r-plus 100.0% 50.0% 100.0% 0.0% 0.0%
mistral-large 100.0% 50.0% 50.0% 0.0% 0.0%
mistral-medium 100.0% 50.0% 50.0% 0.0% 0.0%
mixtral-8x7b 100.0% 25.0% 50.0% 0.0% 0.0%
gpt-3.5-turbo-0125 82.5% 50.0% 95.0% 0.0% 0.0%
gpt-4-0613 100.0% 50.0% 100.0% 0.0% 0.0%
gpt-4-turbo-2024-04-09 92.5% 50.0% 100.0% 0.0% 0.0%
gpt-4o-2024-05-13 95.0% 50.0% 100.0% 0.0% 0.0%
qwen1.5-110B-Chat 100.0% 50.0% 100.0% 0.0% 0.0%
qwen1.5-32B-Chat 95.0% 25.0% 100.0% 0.0% 0.0%
qwen1.5-72B-Chat 100.0% 50.0% 50.0% 0.0% 0.0%
dbrx-instruct 92.5% 50.0% 65.0% 50.0% 50.0%
llama-3-70b-chat 100.0% 50.0% 100.0% 0.0% 0.0%
llama-3-8b-chat 82.5% 50.0% 50.0% 50.0% 50.0%
gemini-1.0-pro 75.0% 25.0% 69.5% 0.0% 0.0%
gemini-1.5-pro 100.0% 50.0% 100.0% 0.0% 0.0%
Table 11: Judging calibration results for 20 LLMs with using granular comparison options without a tie, with $temperature=0$.
Appendix C Further Details on Main Experiment
Figure 8: Distribution of response lengths for 20 LLMs on our EI task, measured in number of tokens.
All votes Consistent votes
LLM Position bias (first) Position bias (second) Self bias Length bias Position bias (first) Position bias (second) Self bias Length bias
qwen1.5-110B-Chat 26.6% 5.8% 0.03 0.44 0.00% $0$04 0.31
gpt-4o-2024-05-13 47.5% 1.7% 0.08 0.45 0.00% 0.00% 0.13 0.22
gpt-4-turbo-2024-04-09 59.0% 2.5% 0.01 0.35 0.00% 0.00% 0.11 0.18
gemini-1.0-pro 5.2% 60.0% -0.01 0.52 0.00% $0$02 0.35
claude-3-opus 9.2% 16.2% -0.08 0.36 0.00% $0$ 0.37
qwen1.5-32B-Chat 75.5% 1.0% 0.00 0.77 0.00% $0$11 0.28
qwen1.5-72B-Chat 0.4% 72.7% 0.00 0.60 0.00% 0.00% 0.07 0.31
llama-3-70b-chat 46.9% 1.9% 0.11 0.50 0.00% 0.00% 0.24 0.3
claude-3-sonnet 4.0% 56.3% 0.11 0.66 0.00% 0.00% 0.24 0.48
dbrx-instruct 52.0% 3.8% 0.03 0.63 0.00% 0.00% 0.04 0.32
claude-3-haiku 52.1% 3.7% 0.07 0.62 0.00% 0.00% 0.17 0.34
command-r-plus 45.1% 2.1% 0.06 0.55 0.00% 0.00% 0.13 0.39
command-r 7.4% 38.1% -0.08 0.62 0.00% $0$1 0.42
mixtral-8x7b 8.2% 33.2% 0.07 0.52 0.00% 0.00% 0.15 0.4
mistral-large 4.4% 23.1% -0.07 0.32 0.00% $0$07 0.21
llama-3-8b-chat 71.7% 2.2% 0.21 0.51 0.00% 0.00% 0.55 0.28
mistral-medium 11.8% 29.2% -0.02 0.53 0.00% 0.00% 0.01 0.41
gpt-4-0613 37.8% 8.6% -0.06 0.42 0.00% $0$04 0.23
gpt-3.5-turbo-0125 32.7% 9.6% 0.04 0.42 0.00% $0$ 0.29
gemini-1.5-pro 1.6% 46.1% 0.14 0.26 0.00% $0$02 0.29
Average Judge 30.0% 20.9% 0.03 0.50 0.0% 0.0% 0.07 0.32
council (by majority vote) 21.5% 3.2% 0.36 3.10% 0.10% 0.35
council (by mean pooling) 26.5% 5.0% 0.59 1.80% 0.90% 0.39
council (no aggregation) 1.6% 46.1% 0.54 0.00% 0.00% 0.35
Table 12: Extended Council judging profiles for bias, with and without consistent votes.
All votes Consistent votes
LLM Contrarianism Agrees most with Disagrees most with Contrarianism Agrees most with Disagrees most with
qwen1.5-110B-Chat 19.2% gpt-4o-2024-05-13 qwen1.5-72B-Chat 8.30% qwen1.5-72B-Chat llama-3-8b-chat
gpt-4o-2024-05-13 18.8% gpt-4-turbo-2024-04-09 qwen1.5-72B-Chat 5.20% gemini-1.5-pro llama-3-8b-chat
gpt-4-turbo-2024-04-09 21.4% gpt-4o-2024-05-13 qwen1.5-72B-Chat 5.90% gemini-1.5-pro llama-3-8b-chat
gemini-1.0-pro 43.3% qwen1.5-72B-Chat qwen1.5-32B-Chat 17.90% gpt-4o-2024-05-13 llama-3-8b-chat
claude-3-opus 20.6% mistral-large llama-3-8b-chat 13.80% qwen1.5-72B-Chat llama-3-8b-chat
qwen1.5-32B-Chat 32.2% llama-3-8b-chat qwen1.5-72B-Chat 9.70% gpt-4-turbo-2024-04-09 llama-3-8b-chat
qwen1.5-72B-Chat 46.6% claude-3-sonnet qwen1.5-32B-Chat 7.80% qwen1.5-110B-Chat llama-3-8b-chat
llama-3-70b-chat 22.3% gpt-4-turbo-2024-04-09 qwen1.5-72B-Chat 8.20% gemini-1.5-pro gemini-1.0-pro
claude-3-sonnet 40.1% qwen1.5-72B-Chat qwen1.5-32B-Chat 13.10% gpt-4-turbo-2024-04-09 command-r
dbrx-instruct 24.5% gpt-4-turbo-2024-04-09 qwen1.5-72B-Chat 9.50% gpt-4o-2024-05-13 llama-3-8b-chat
claude-3-haiku 27.6% llama-3-70b-chat qwen1.5-72B-Chat 13.00% llama-3-70b-chat qwen1.5-32B-Chat
command-r-plus 22.8% gpt-4-turbo-2024-04-09 qwen1.5-72B-Chat 8.80% gemini-1.5-pro llama-3-8b-chat
command-r 33.5% gemini-1.5-pro llama-3-8b-chat 15.30% gpt-4-turbo-2024-04-09 llama-3-8b-chat
mixtral-8x7b 33.5% gemini-1.5-pro llama-3-8b-chat 15.90% qwen1.5-72B-Chat gemini-1.0-pro
mistral-large 21.2% claude-3-opus llama-3-8b-chat 6.00% gemini-1.5-pro llama-3-8b-chat
llama-3-8b-chat 36.0% qwen1.5-32B-Chat qwen1.5-72B-Chat 25.70% llama-3-70b-chat gpt-4-turbo-2024-04-09
mistral-medium 30.5% mistral-large llama-3-8b-chat 12.20% qwen1.5-72B-Chat llama-3-8b-chat
gpt-4-0613 20.3% gpt-4-turbo-2024-04-09 qwen1.5-72B-Chat 7.90% gpt-4o-2024-05-13 llama-3-8b-chat
gpt-3.5-turbo-0125 25.1% gpt-4o-2024-05-13 qwen1.5-72B-Chat 12.80% gpt-4o-2024-05-13 llama-3-8b-chat
gemini-1.5-pro 33.2% mistral-large llama-3-8b-chat 4.00% gpt-4-turbo-2024-04-09 llama-3-8b-chat
Average Judge 28.6% 11.1%
council (by majority vote) gpt-4o-2024-05-13 qwen1.5-72B-Chat gemini-1.5-pro llama-3-8b-chat
council (by mean pooling) mistral-large qwen1.5-72B-Chat 0.10% gemini-1.5-pro llama-3-8b-chat
council (no aggregation)
Table 13: Extended Council judging profiles for agreement, with and without consistent votes.
All votes Consistent votes
LLM Polarization Lowest affinity for Highest affinity for Polarization Lowest affinity for Highest affinity for
qwen1.5-110B-Chat 62.6 gemini-1.5-pro qwen1.5-110B-Chat 78.30% gemini-1.5-pro qwen1.5-110B-Chat
gpt-4o-2024-05-13 65.4 gemini-1.5-pro gpt-4o-2024-05-13 84.80% gemini-1.5-pro gpt-4o-2024-05-13
gpt-4-turbo-2024-04-09 54.5 gpt-3.5-turbo-0125 gpt-4o-2024-05-13 88.20% mistral-medium gpt-4o-2024-05-13
gemini-1.0-pro 31.0 gemini-1.5-pro qwen1.5-110B-Chat 72.30% gemini-1.5-pro qwen1.5-110B-Chat
claude-3-opus 73.0 gpt-3.5-turbo-0125 qwen1.5-110B-Chat 93.30% gemini-1.5-pro qwen1.5-110B-Chat
qwen1.5-32B-Chat 46.7 gemini-1.5-pro gpt-4-turbo-2024-04-09 87.50% gpt-3.5-turbo-0125 qwen1.5-110B-Chat
qwen1.5-72B-Chat 45.7 gemini-1.5-pro qwen1.5-110B-Chat 84.90% gemini-1.5-pro qwen1.5-110B-Chat
llama-3-70b-chat 68.3 gemini-1.5-pro qwen1.5-110B-Chat 89.30% gemini-1.5-pro qwen1.5-110B-Chat
claude-3-sonnet 49.7 gemini-1.5-pro gpt-4o-2024-05-13 82.60% gemini-1.5-pro gpt-4o-2024-05-13
dbrx-instruct 54.5 gemini-1.5-pro qwen1.5-110B-Chat 85.00% gemini-1.5-pro qwen1.5-110B-Chat
claude-3-haiku 51.1 gemini-1.5-pro qwen1.5-110B-Chat 79.40% gemini-1.5-pro qwen1.5-110B-Chat
command-r-plus 52.5 gemini-1.5-pro qwen1.5-110B-Chat 78.40% gemini-1.5-pro qwen1.5-110B-Chat
command-r 44.4 gemini-1.5-pro qwen1.5-110B-Chat 53.80% gemini-1.5-pro gemini-1.0-pro
mixtral-8x7b 59.4 gemini-1.5-pro qwen1.5-110B-Chat 81.30% gemini-1.5-pro qwen1.5-110B-Chat
mistral-large 78.8 gemini-1.5-pro qwen1.5-110B-Chat 92.00% gpt-3.5-turbo-0125 qwen1.5-110B-Chat
llama-3-8b-chat 34.9 gpt-3.5-turbo-0125 llama-3-70b-chat 76.20% gpt-3.5-turbo-0125 llama-3-70b-chat
mistral-medium 58.0 gemini-1.5-pro qwen1.5-110B-Chat 81.90% gemini-1.5-pro qwen1.5-110B-Chat
gpt-4-0613 62.0 gemini-1.5-pro qwen1.5-110B-Chat 86.00% gemini-1.5-pro qwen1.5-110B-Chat
gpt-3.5-turbo-0125 65.6 gemini-1.5-pro qwen1.5-110B-Chat 84.20% gemini-1.5-pro qwen1.5-110B-Chat
gemini-1.5-pro 61.7 gpt-3.5-turbo-0125 qwen1.5-110B-Chat 90.00% gemini-1.5-pro gpt-4o-2024-05-13
Average Judge 56.0 82.47%
council (by majority vote) 77.0 gemini-1.5-pro qwen1.5-110B-Chat 82.50% gemini-1.5-pro qwen1.5-110B-Chat
council (by mean pooling) 60.3 gemini-1.5-pro qwen1.5-110B-Chat 80.50% gemini-1.5-pro qwen1.5-110B-Chat
council (no aggregation) 54.0 gemini-1.5-pro qwen1.5-110B-Chat 81.10% gemini-1.5-pro qwen1.5-110B-Chat
Table 14: Extended Council judging profiles for affinity, with and without consistent votes.
Figure 9: Full heatmap of LLM judge to LLM respondent affinities from the main experiment.
Figure 10: Full heatmap of LLM judge to LLM respondent affinities from the main experiment, with Council consensus affinity subtracted out.
Figure 11: Graph of top 5 affinities. An edge exists from LLM $a$ to LLM $b$ if $\text{affinity}(a,b)$ is in the top 5 affinities for LLM $a$.
Figure 12: Full heatmap of LLM judge to LLM judge Cohen’s Kappa sidewise agreement scores from the main experiment.
Figure 13: Graph of top 5 agreement. An edge exists from LLM $a$ to LLM $b$ if $\text{agreement}(a,b)$ is in the top 5 agreement for LLM $a$.
Figure 14: Expected LLM respondent vs. LLM respondent win rates derived from Terry-Bradley coefficients.
Appendix D Human Evaluation
During registration for our experiments, all candidates provided their demographic details (see Figure 16. Additionally, we required each candidate to complete a questionnaire measuring their level of empathy, sourced from [21]. All candidates were informed of the purpose of our study. 142 participants completed the survey but after removing those who failed attention checks, 102 participants remain. Each dilemma pair and response was rated by 11 participants on average, after removing malicious participants. Each participant was compensated £9.00 per hour.
Participant demographics:
All participants are over 18 years old. Our sample is made up of 53 women, 46 men, and one non-binary identifying individual. 84 of our participants were from the United Kingdom, 14 from the United States and two from other English-speaking countries; all were native English speakers. With regards to their use of AI chatbots, 23 report using them every day or nearly every day, 48 sometimes, four rarely and only four report never using them. None report having difficulties reading long texts. None report having difficulties reading long texts.
Data quality assurance:
Because the task is both difficult and subjective, we take a two-fold approach to ensure quality data: (1) we ask participants to provide demographics which we cross-reference with data from Prolific; and (2) we use two repeated dilemmas as test questions, checking for self-agreement. We allow participants to shift slightly to account for the lack of ties: a participant may slightly prefer one response then another, but not prefer one strongly then prefer a different response the following time. We remove data from workers who lack this consistency. This results in 102 unique participants in the final set.
We provide the participant guidelines in Figures 17 and 18.
Measuring perceived empathy:
We adapt our feedback from the scale proposed by [34], which is designed to assess systems with which the users have interacted. We exclude question E5 from the original questionnaire and rephrase them to fit our experiment. The statements are detailed in Table 15.
E1 The best response considered the protagonist’s mental state.
E2 (EQ) The best response seemed emotionally intelligent.
E3 The best response expressed emotions.
E4 The best response sympathized with the protagonist.
E5 The best response was supportive in coping with an emotional situation.
U1 The best response understood the protagonist’s goals.
U2 The best response understood the protagonist’s needs.
U3 The best response seems trustworthy.
U4 The best response understood the protagonist’s intentions.
Table 15: Adapted PETS scale for our study.
Figure 15: Proportion of times users found the statements in the PETS questionnaire to be true about the winning response. The corresponding statements are shown in Table 15. E2 in the questionnaire is equivalent to out EQ question (shown first) so it is not included.
Figure 16: Participant demographic questionnaire.
Figure 17: Participant guidelines for rating the generation of dilemmas.
Figure 18: Participant guidelines for rating the responses to dilemmas.
Appendix E Further Details on Qualitative Analysis of Council Judgments
We provide full figures for all 38 coarse-grained reasons that the council of LLM judges uses to justify their preferences.
Figure 19: Citation frequency of 38 qualitative reasons why the winning response was preferred.
Figure 20: Spearman correlation matrix of cited reasons why the winning response was preferred.
Appendix F Prompt Templates
In this section, we list all prompts used, including prompts for synthetic expansion, dilemma response, and judging.
Figure 21: Prompt used to convert EmoBench Emotional Application data into scenarios with richer detail.
Figure 22: Prompt for primary emotional application task: respond to a nuanced emotional interpersonal dilemma.
Figure 23: Prompt used to assess whether an expanded scenario would be appropriate to include in an emotional intelligence test.
Figure 24: Prompt used for pairwise comparison between responses.
Figure 25: Prompt variations on Figure 24 (applied to the bottom highlighted text) used to study natural consistency and variability under different pairwise comparison regimes in Appendix B.
Figure 26: Prompt used to map explanations in pairwise ratings to a rich, fixed set of qualitative reasons. The 38 seed qualitative reasons used in the prompt come from manual review of 50 randomly selected pairwise ratings in the main experiment involving the full council of 20 LLMs.
Appendix G Datasheet
We follow documentation practices described in Datasheets for Datasets  [^6] https://arxiv.org/abs/1803.09010.
Motivation
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
LMC-EA was developed to demonstrate how to benchmark foundation models on highly subjective tasks such as those in the domain of emotional intelligence by the collective consensus of a council of LLMs.
Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
This dataset was created by the authors of this paper.
Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
Predibase
Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
There are 4 parts of LMC-EA dataset:
  1. 1.
    Test set formulation: Synthetic expansions of the EmoBench EA dataset [^7] https://github.com/Sahandfer/EmoBench/blob/master/data/EA/data.json, generated by 20 different LLMs. Each expansion is a detailed story describing an interpersonal conflict, written in the first person.
  2. 2.
    Response collection: Conversational responses to 100 interpersonal conflicts, from 20 different LLMs. The prompt to an LLM for a conversational response requests that the response is at most 250 words in response length.
  3. 3.
    Response judging (council): LLM ratings for pairwise comparisons for every non-reference LLM’s response vs. the reference LLM’s response, for each interpersonal conflict, from each LLM judge. To mitigate position bias, we adopt a two-game setup, swapping model positions per query.
  4. 4.
    Response judging (human): Ratings for pairwise comparisons for a subset of 9 LLMs and 120 randomly sampled dilemma-response tuples. We recruited a total of 142 participants.
How many instances are there in total (of each type, if appropriate)?
  1. 1.
    Test set formulation: There are 200 interpersonal conflicts.
  2. 2.
    Response collection: There are 100 interpersonal conflicts x 20 LLMs = 2000 responses.
  3. 3.
    Response judging (council): There are 100 interpersonal conflicts x 19 non-reference LLM responses x 20 LLM judges x 2 position swaps = 76000 responses.
  4. 4.
    Response judging (human): Each dilemma response pair was rated by 11 participants on average, with a total of 1343 ratings.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
Due to budget constraints, response collection and response judging is performed on a subset of 100 interpersonal conflicts out of the full set of 200 interpersonal conflicts from the original EmoBench dataset. The 100 interpersonal conflicts is representative of a diverse set of interpersonal problems.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images) or features? In either case, please provide a description.
See main paper or the dataset link for examples.
Is there a label or target associated with each instance? If so, please provide a description.
No.
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
No.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
No, except for the emobench_id across subsets can be used to trace a full path from original EmoBench scenario → synthetic expansion → conversational response → response judging.
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
The LMC-EA dataset is expected to be used only for testing purposes.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
The extraction of the exact pairwise rating (A>>B, A>B, B>A, B>>A) in response judging is performed by regular expressions and other heuristics-based substring presence rules. Although we manually checked and assigned responses for which an exact pairwise rating could not be automatically extracted, there might be corner error cases that may have been missed.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? If it links to or relies on external resources, a) are there guarantees that they will exist, and remain constant, over time; b) are there official archival versions of the complete dataset (i.e., including the external resources as they existed at the time the dataset was created); c) are there any restrictions (e.g., licenses, fees) associated with any of the external resources that might apply to a future user? Please provide descriptions of all external resources and any restrictions associated with them, as well as links or other access points, as appropriate.
The data is self-contained.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals non-public communications)? If so, please provide a description.
No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
No, to the best of our knowledge.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
Our dataset is composed of hypothetical scenarios designed to simulate various conflict situations. These scenarios are entirely fictional and have been crafted for the purpose of research and analysis. Any resemblance to actual persons, living or dead, is purely coincidental.
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
No.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
No.
Does the dataset contain data that might be considered sensitive in any way (e.g., data that reveals racial or ethnic origins, sexual orientations, religious beliefs, political opinions or union memberships, or locations; financial or health data; biometric or genetic data; forms of government identification, such as social security numbers; criminal history)? If so, please provide a description.
No, to the best of our knowledge.
Collection Process
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
Responses from LLMs were generated by open source and proprietary LLMs, using carefully designed prompts.
For human ratings, we recruit participants via crowdsourcing on Prolific [^8] https://www.prolific.com/.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
LLM outputs were obtained through a variety of providers and APIs (Table 16). For conversational response collection, the API’s default temperature was used. For response judging, a temperature of 0 was used.
Organization LLM Provider and API
Open AI gpt-4o-2024-05-13 OpenAI API (https://platform.openai.com/docs/api-reference)
Open AI gpt-4-turbo-04-09 OpenAI API (https://platform.openai.com/docs/api-reference)
Open AI gpt-4-0613 OpenAI API (https://platform.openai.com/docs/api-reference)
Open AI gpt-3.5-turbo-0125 OpenAI API (https://platform.openai.com/docs/api-reference)
Mistral mistral-large-latest Mistral AI API (https://docs.mistral.ai/api/)
Mistral open-mixtral-8x22b Mistral AI API (https://docs.mistral.ai/api/)
Mistral open-mixtral-8x7b Mistral AI API (https://docs.mistral.ai/api/)
Meta llama-3-70b-chat-hf Together REST API (https://docs.together.ai/docs/inference-rest)
Meta llama-3-8b-chat-hf Together REST API (https://docs.together.ai/docs/inference-rest)
Google gemini-1.5-pro-preview-0409 Vertex AI API (https://cloud.google.com/vertex-ai/docs/reference/rest)
Google gemini-1.0-pro Vertex AI API (https://cloud.google.com/vertex-ai/docs/reference/rest)
Databricks dbrx Together REST API (https://docs.together.ai/docs/inference-rest)
Cohere command-r-plus Cohere API (https://docs.cohere.com/reference/chat)
Cohere command-r Cohere API (https://docs.cohere.com/reference/chat)
Anthropic claude-3-opus-20240229 Anthropic API (https://docs.anthropic.com/en/api/messages)
Anthropic claude-3-sonnet-20240229 Anthropic API (https://docs.anthropic.com/en/api/messages)
Anthropic claude-3-haiku-20240307 Anthropic API (https://docs.anthropic.com/en/api/messages)
Alibaba qwen1.5-110B-chat Together REST API (https://docs.together.ai/docs/inference-rest)
Alibaba qwen1.5-72B-chat Together REST API (https://docs.together.ai/docs/inference-rest)
Alibaba qwen1.5-32B-chat Together REST API (https://docs.together.ai/docs/inference-rest)
Table 16: List of Language Model Council LLMs and providers and APIs used.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
EmoBench scenarios ids 100-199 are used.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
LLM responses were collected by the authors with APIs listed above.
For the human study on response judging, all participants are over 18 years old. Our sample is made up of 53 women, 46 men, and one non-binary identifying individual. 84 of our participants were from the United Kingdom, 14 from the United States and two from other English-speaking countries; all were native English speakers. With regards to their use of AI chatbots, 23 report using them every day or nearly every day, 48 sometimes, four rarely and only four report never using them. None report having difficulties reading long texts.
We have a total of 102 participants. Each dilemma pair and response was rated by 11 participants on average, after removing malicious participants. Each participant was compensated £9.00 per hour.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
The dataset was collected in April and May of 2024.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
No.
Does the dataset relate to people? If not, you may skip the remaining questions in this section.
No.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
For human ratings, participants are recruited through Prolific [^9] https://www.prolific.com/.
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
No.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
Yes.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
Yes, Prolific allows workers to revoke consent.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
N/A.
Preprocessing/cleaning/labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
No.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
N/A.
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
N/A.
Uses
Has the dataset been used for any tasks already? If so, please provide a description.
Yes, for experiments described in the main paper.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
https://huggingface.co/datasets/llm-council/emotional_application
What (other) tasks could the dataset be used for?
The dataset is designed to test the ability of a council of LLMs to evaluate each other in a full consensus manner.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks) If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
No.
Are there tasks for which the dataset should not be used? If so, please provide a description.
No.
Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
Yes.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub) Does the dataset have a digital object identifier (DOI)?
The dataset is publicly available through the https://huggingface.co/datasets/llm-council/emotional_application, which supports direct download or loading the dataset in Python [^10] https://huggingface.co/docs/datasets/en/loading.
When will the dataset be distributed?
The dataset is distributed in June 2024.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
Yes, CC-BY [^11] https://creativecommons.org/licenses/by/4.0/ license.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
No, to the best of our knowledge.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No, to the best of our knowledge.
Maintenance
Who will be supporting/hosting/maintaining the dataset?
The authors of this publication.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
Yes, by email or any other contact point provided at the top of this document.
Is there an erratum? If so, please provide a link or other access point.
No.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
No updates are planned at the moment. If any is made, it will be communicated at https://huggingface.co/datasets/llm-council/emotional_application.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
N/A.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
Yes.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
Please contact the dataset maintainers using the contact information above or start a discussion at https://huggingface.co/datasets/llm-council/emotional_application.